# X-ray Variability in the Symbiotic Binary RT Cru: Principal Component Analysis

### v1

@@ Abstract
Hard X-ray-emitting (-type) symbiotic binaries, which exhibit a strong hard X-ray excess, have posed a challenge to our understanding of accretion physics in degenerate dwarfs. RT Cru, which is a member of -type symbiotics, shows stochastic X-ray variability. Timing analyses of X-ray observations from XMM-Newton and NuSTAR, which we consider here, indicate hourly fluctuations, in addition to a spectral transition from 2007 to a harder state in 2012 seen with Suzaku observations. To trace the nature of X-ray variability, we analyze the multi-mission X-ray data using principal component analysis (PCA), which determines the spectral components mostly contributing to the flickering behavior and the hardness transition. The Chandra HRC-S/LETG and XMM-Newton EPIC-pn data provide the primary PCA components, which may contain some variable emission features, especially in the soft excess. Additionally, the absorbing column (first order with 50%), along with the source continuum (20%), and a third component (9%) – which likely accounts for thermal emission in the soft band – are the three principal components found in the Suzaku XIS1 observations. The PCA components of the NuSTAR data also correspond to the continuum and possibly emission features. Our findings suggest that the spectral hardness transition between the two Suzaku observations is mainly due to changes in the absorbing material and X-ray continuum, while some changes in the thermal plasma emission may result in flickering-type variations.
]Received 2024 December 31; revised 2024 June 22; accepted 2024 June 23
@@ 1 Introduction
Symbiotic systems refer to binary stars that are characterized by the presence of a hot degenerate core accreting matter from a red giant cool star (Paczynski & Rudak, 1980; Kenyon & Webbink, 1984; Belczyński et al., 2000). They exhibit soft or supersoft thermal X-ray emission (Muerset et al., 1997; Luna et al., 2013) dominated by the blackbody-like or bremsstrahlung radiation (Imamura & Durisen, 1983). However, a small group of them have been observed to emit extreme hard X-ray excess above 2.4 keV (Tueller et al., 2005a; Bird et al., 2007; Kennea et al., 2009; Eze, 2014). These are the hard X-ray emitting symbiotics or -type sources according to the classification scheme by Luna et al. (2013). They could be progenitors of type Ia supernovae owing to the possible presence of massive white dwarfs (Kennea et al., 2009). Among this group, notable systems can be mentioned: RT Cru (Luna & Sokoloski, 2007; Ducci et al., 2016; Luna et al., 2018; Danehkar et al., 2021), CH Cyg (Wheatley & Kallman, 2006; Mukai et al., 2007; Toalá et al., 2023), T CrB (Luna et al., 2008; Zhekov & Tomov, 2019), SS73 17 (CD–57 3057; Smith et al., 2008; Eze et al., 2010), and MWC 560 (Stute & Sahai, 2009; Lucy et al., 2020). The discovery of this particular group presents a challenge to our knowledge about accretion physics in white dwarfs due to the strong hard X-ray emission (Luna & Sokoloski, 2007; Kennea et al., 2009). Recently, Toalá (2024) proposed a disc-like model to explain the X-ray properties of symbiotics, where -type sources have an accretion disk near the edge. In addition, radiative transfer simulations of X-ray photons by Toalá (2024) associated the -type group with a low-accreting degenerate core with high-temperature plasma ( keV) within the boundary layer between the inner edge of the accretion disk and the white dwarf surface.
@@
Some of hard X-ray-emitting symbiotics seem to produce distinctive soft and hard thermal plasma emissions: CH Cyg with plasma temperatures of 0.2, 0.7, and 7.3 keV (Ezuka et al., 1998), SS73 17 with and keV temperatures (Eze et al., 2010), and MWC 560 showing thermal emissions with temperatures of and keV (Stute & Sahai, 2009). Symbiotic stars with distinctive soft and hard X-ray thermal components are referred to as the -type group (Luna et al., 2013). Previous studies have also identified the presence of jets in some of them, namely CH Cyg (Galloway & Sokoloski, 2004; Karovska et al., 2007, 2010) and MWC 560 (Tomov et al., 1992; Schmid et al., 2001; Lucy et al., 2018). The soft thermal emission found in these system might have a potential link to the shock region created by either a jet or interaction of a wind colliding with the surrounding material (Stute & Sahai, 2009). According to Toalá (2024), the soft X-ray emission of a two-temperature plasma model can be obscured by the disk, so extended emission from jets, colliding winds, and/or hot bubbles is likely responsible for the soft component in sources. The soft X-ray component could originate from colliding stellar winds (CSW), though a different mechanism such as accretion may also be responsible for the X-ray emission seen in some epochs in the -class symbiotic star AG Peg (Zhekov & Tomov, 2016).
@@
The X-ray variability of RT Cru recorded with the Chandra telescope has been recently investigated using hardness ratio analysis and spectral modeling (Danehkar et al., 2021). To further evaluate the nature of the X-ray variability in this -type symbiotic star and the soft thermal plasma emission suggested by Danehkar et al. (2021), we conduct further timing analyses on the archival data taken with XMM-Newton, Suzaku, and NuSTAR, in addition to comprehensive eigenvector-based multivariate analyses of the historical X-ray data of the Chandra and other telescopes. Section 2 describes our data reduction of time-sliced data required for implementing principal component analysis (PCA). In Section 3, we investigate the X-ray light curves and hardness ratios. Section 4 presents our principal components determined with PCA, as well as simulated PCA spectra, followed by discussions in Section 5 and a conclusion in Section 6.
@@ 2 Data Reduction for PCA
RT Cru was observed using the X-ray Imaging Spectrometer (XIS; Koyama et al., 2007) and the Hard X-ray Detector (HXD; Takahashi et al., 2007) on board Suzaku (Mitsuda et al., 2007) in 2007 and 2012, the EPIC-pn instrument (Strüder et al., 2001) aboard the XMM-Newton telescope (Jansen et al., 2001) in 2019, the ‘A’ and ‘B’ focal plane modules (FPM) on the Nuclear Spectroscopic Telescope Array (NuSTAR) satellite (Harrison et al., 2013) in 2016, and the Low Energy Transmission Grating (LETG; Brinkman et al., 2000) on the High Resolution Camera Spectrometer (HRC-S; Murray et al., 2000) of Chandra X-ray Observatory (CXO; Weisskopf et al., 2000, 2002) in 2015. The observations are summarized in Table 1, which includes the instrument and its configuration, observation identification number, start and end time (UTC), exposure time (ks), and total counts and count-rate (count s-1) of each observation. To implement eigenvector-based multivariate statistical analysis such as PCA of a variable X-ray source, it is necessary to split each dataset into a time series of spectral data at a fixed interval of 10 ks.
@@
We downloaded the Chandra HRC-S/LETG event data from the Chandra data archive and reprocessed them using the chandra_repro tool from the ciao package (version 4.15; Fruscione et al., 2006) together with the Chandra CALDB data (version 4.10.2).111The Chandra dataset is contained in https://doi.org/10.25574/cdc.201 (catalog doi:10.25574/cdc.201). The time-segmented event files were produced by applying the ciao operations dmcopy and dmappend on the 2nd level events. The spectra, together with their respective redistribution and response data, were generated by applying the ciao programs tgextract and mktgresp to the time-sliced event files. The dmtype2split tool, part of the ciao software, was utilized to segregate distinct positive and negative orders, while the application tg_bkg was used to create the background spectra.
@@
The XMM-Newton EPIC-pn data were obtained from XMM Science Archive and reduced with the science analysis software (sas v 20.0.0; Gabriel et al., 2004) and the calibration files (XMM-CCF-REL-391). The use of the sas tool epproc led to the generation of processed event files. These events were subsequently employed to produce new event files, which were stacked at an interval of 10 ks with the sas program evselect. We removed the events of flaring particle background, which were identified using the count rates of c s-1 in the single-pixel (PATTERN 0) light curves binned at 100 sec within the 10–12 keV energy range. Furthermore, the time-filtered event files exclusively included single and double patterned events (PATTERN 4) within the relevant PI channel range (200 PI 15000), disregarding defective pixels (FLAG 0). The procedure especget was utilized to generate a set of source spectra by applying it to the time-sliced event data. The spectra of the source were taken from a circular region with a radius of centered on the brightest peak of the source. The background spectra were created using a circle of the same size on the same chip, but without any sources included.
@@
The Suzaku data acquired using the XIS and HXD-PIN were retrieved and processed using the ftool program aepipeline from the HEAsoft software (v 6.31.1) and the calibration data (XIS: 2018-10-10 and HXD: 2011-09-13). The XIS provided energy coverage ranging from 0.2 to 12 keV, while the HXD supplied a band pass of 10-70 keV using PIN diodes. Various Good Time Interval (GTI) files at our chosen time-interval were created using the ftool application maketime. The GTI files were employed in the multipurpose tool xselect to generate a set of time-segmented source spectra. The source spectrum was acquired by extracting data from a circular region with a radius of centered on the emission peak of the source, whereas the background was chosen from a nearby circular region of the same radius excluding any sources. We should specify that XIS1 was the only back-illuminated (BI) detector aboard Suzaku, whereas the other XIS devices were front-illuminated (FI). Typically, BI detectors are expected to offer superior quantum efficiency in comparison to FI detectors under an identical depletion depth (Lesser & Iyer, 1998). The program hxdpinxbpi was employed to generate the time-sliced spectra of HXD-PIN data using the GTI files and the ”tuned” non-X-ray PIN background released by the Suzaku team.222https://heasarc.gsfc.nasa.gov/FTP/suzaku/data/background/
@@
The NuSTAR data taken with the FPM ‘A’ and ‘B’ were downloaded from the HEASARC archive and reduced using the tool nupipeline from NuSTARDAS (v 2.1.2) and the relevant calibration files (8 Feb 2023). The ftool program maketime was used to create the GTI tables at an interval of 10 ks, which were then applied to the calibrated event files via the application nuproducts resulting in the time-stacked spectra. The source was extracted it from a circular aperture with a radius of , whereas the background was from a location on the same chip that was devoid of any sources.
@@ 3 Time Series Analysis
The first two equations are identical to those employed by Prestwich et al. (2003) for the classification of X-ray sources in the Local Group galaxies. The hardness ratio diagrams are created by plotting the hardness ratios against the entire bands, which assist in the detection of spectral transitions associated with accretion or blockage made by absorbing material. Hardness ratio analysis has been employed to characterize various astronomical objects, including extragalactic X-ray sources (e.g., Hong et al., 2004; Plucinsky et al., 2008), quasars (Danehkar et al., 2018; Boissay-Malaquin et al., 2019), and X-ray binaries (Sreehari & Nandi, 2021).
In Figure 1, we present the background-subtracted light curves of RT Cru made with the XMM-Newton, Suzaku, and NuSTAR observations binned at intervals of 600 sec in the broad (; 0.4–10 keV), hard (), and extreme hard () bands. The figure also shows the corresponding hardness ratios, namely and for XMM-Newton (pn) and Suzaku (XIS1), and for NuSTAR and Suzaku (XIS1+PIN). The uncertainties in the light curves and hardness ratios were determined using the Bayesian Estimator for Hardness Ratios (BEHR; Park et al., 2006). Hour-scale variations are evident in all the time series. In particular, the source got harder in February 2012 compared to July 2007 according to the time series of the Suzaku XIS1 data, whereas the source brightness () in 2012 is lower than that in 2007. Moreover, the and hardness ratios involving the extreme hard band () were stronger in February 2012.
Figure 2 shows hardness ratio diagrams that illustrate the relationship between the hardness ratios and the broad bands ( and ). The hardness diagrams of the XMM-Newton observation do not depict a hardness transition similar to what was seen with the Chandra data (Danehkar et al., 2021). However, the and diagrams of the two Suzaku observations indicate that the source experienced a transition from the high/soft to low/hard spectral states between 2007 and 2012. Moreover, the Suzaku diagrams also revealed that the source was stronger in the soft (0.4–1.1 keV) and hard (2.6–10 keV) bands – lower and higher at the same time – in February 2012 by comparison with those in July 2007. Although we do not see any special pattern in the hardness ratio diagrams of NuSTAR, the and diagrams of the Suzaku light curves imply a higher average of the extreme hardness () in February 2012.
@@ 3.1 Statistical Tests of X-ray Variability
where and are the standard deviation and the mean of data points, respectively, is the index of each data point in the time series, and is the total number of data points. This coefficient describes the extent of variations with respect to the mean, so its higher values are associated with stronger variability if we compare the same parameter. However, this coefficient is not suitable for comparison between dissimilar parameters and cannot evaluate the intrinsic nature of variations. The ratio of the standard deviation of errors to that of data points () may help us see whether the variability is intrinsic to the observation. Again, this ratio is suitable for evaluating values for the identical parameter. Both and are not able to characterize detailed features of consecutive series. The normalized consecutive number, so-called Con, which was first used by Wozniak (2000), may be used to quantify the changes in a number of consecutive points. To evaluate continuity in light curves, we define the Con as the number of three consecutive points greater or lower than normalized by . For random fluctuations, Con has a value of and is equal to 0 for constant consecutive series.
@@
For a normal distribution, the mean von Neumann ratio is expected to be (Young, 1941). To assess the presence of autocorrelation, we estimated the confidence intervals of with the significance level of for a normal probability distribution. The von Neumann ratio within the confidence levels of implies no autocorrelation, whereas values outside the aforementioned confidence levels toward 0 and 4 correspond to positive and negative autocorrelation in successive data points, respectively.
@@
Table 2 summarizes the results of our statistical analysis of the X-ray variability in RT Cru obtained with different methods. It can be seen that and of , HR1, and HR2, are higher in the XMM-Newton data than those in the Suzaku observations. This implies that variations are higher in the XMM-Newton light curves, but with larger uncertainties. We also notice a higher in the HR4 ratio of the NuSTAR data, which may be an indication of scattered fluctuations in this ratio. However, is not a suitable tool for quantifying the variability in a time sequence of data. The Con number may be able to better distinguish variations in a consecutive sequence. The values of Con indicate consecutive changes in in the XMM-Newton and combined Suzaku/XIS1 observations, HR2 in the first Suzaku/XIS1 observation and combined two Suzaku/XIS1 observations, while others with Con may have random variations. Nevertheless, our Con statistical method is based only on three consecutive data points, so it is unable to obtain a broader picture of variability in the entire sequence of data points.
@@
In the case of Suzaku/XIS1, the von Neumann ratio demonstrates positive autocorrelation in the time series of the broad band () and hardness ratios of the first, second, and combined observations, apart from HR1 in the second observation. However, the -values of the normality tests suggest that there are not normal distributions in HR1 of the first XIS1 observation with marginally significant statistics and the second XIS1 observation with significant statistics. Furthermore, the A–D and S–W tests of , HR1, and HR2 in the mixed two Suzaku/XIS1 observations definitely reject the hypothesis of normality, which is obvious in the movement pattern seen in the hardness diagrams in Fig. 2. Nevertheless, we see a non-normal distribution only in the HR3 ratio of the Suzaku/PIN observations, but not and HR4. The von Neumann statistical analysis of the Suzaku/PIN data also depicts positive autocorrelation on the broad band of the first and combined observations, and on HR3 and HR4 of the mixed multi-epoch data, which are consistent with the spectral transition occurring between 2007 and 2012.
The broad band () and the hardness ratios (HR3 and HR4) of the NuSTAR data exhibit non-random distributions according to the -values obtained from A–D and S–W tests. However, von Neumann’s mean squared successive difference depicts a positive autocorrelation only in the broad band. The hardness diagrams in Fig. 2 seem to not depict any obvious hardness transition in the NuSTAR data, though our A–D and S–W normality tests apparently suggest it, which is not supported by the von Neumann statistics. However, all of the normality tests and the von Neumann ratio indicate that there are statistically significant fluctuations in the NuSTAR board band.
In summary, our statistical analysis of the variability shows that RT Cru underwent flux variations over the broad-band of the XMM-Newton data based on and the A–D/S–W test, the first and second Suzaku/XIS1 data, the first and combined Suzaku/PIN data, the NuSTAR data according to the von Neumann results, and the mixed first-second Suzaku/XIS1 data and the NuSTAR data based on the A–D/S–W test. Moreover, our results suggest that the X-ray source experienced a long-term spectral transition between the first and second Suzaku observations based on the normality tests of HR1, HR2, and HR3, as well as some short-term hardness fluctuations in HR1 of the second Suzaku/XIS1 data and HR3 of the NuSTAR data.

### v2

@@ Abstract
Hard X-ray-emitting (-type) symbiotic binaries, which exhibit a strong hard X-ray excess, have posed a challenge to our understanding of accretion physics in degenerate dwarfs. RT Cru, which is a member of the -type symbiotics, shows stochastic X-ray variability. Timing analyses of X-ray observations from XMM-Newton and NuSTAR, which we consider here, indicate hourly fluctuations, in addition to a spectral transition from 2007 to a harder state in 2012 seen with Suzaku observations. To trace the nature of X-ray variability, we analyze the multi-mission X-ray data using principal component analysis (PCA), which determines the spectral components that contribute most to the flickering behavior and the hardness transition. The Chandra HRC-S/LETG and XMM-Newton EPIC-pn data provide the primary PCA components, which may contain some variable emission features, especially in the soft excess. Additionally, the absorbing column (first order with 50%), along with the source continuum (20%), and a third component (9%) – which likely accounts for thermal emission in the soft band – are the three principal components found in the Suzaku XIS1 observations. The PCA components of the NuSTAR data also correspond to the continuum and possibly emission features. Our findings suggest that the spectral hardness transition between the two Suzaku observations is mainly due to changes in the absorbing material and X-ray continuum, while some changes in the thermal plasma emission may result in flickering-type variations.
]Received 2023 December 31; revised 2024 June 22; accepted 2024 June 23; published 2024 August 28
@@ 1 Introduction
Symbiotic systems refer to binary stars that are characterized by the presence of a hot degenerate core accreting matter from a cool red giant star (Paczynski & Rudak, 1980; Kenyon & Webbink, 1984; Belczyński et al., 2000). They exhibit soft or supersoft thermal X-ray emission (Muerset et al., 1997; Luna et al., 2013) dominated by blackbody-like or bremsstrahlung radiation (Imamura & Durisen, 1983). However, a small group of them have been observed to emit an extreme hard X-ray excess above 2.4 keV (Tueller et al., 2005a; Bird et al., 2007; Kennea et al., 2009; Eze, 2014). These are the hard X-ray-emitting symbiotics or -type sources according to the classification scheme by Luna et al. (2013). They could be progenitors of type Ia supernovae owing to the possible presence of massive white dwarfs (Kennea et al., 2009). Among this group, notable systems can be mentioned: RT Cru (Luna & Sokoloski, 2007; Ducci et al., 2016; Luna et al., 2018; Danehkar et al., 2021), CH Cyg (Wheatley & Kallman, 2006; Mukai et al., 2007; Toalá et al., 2023), T CrB (Luna et al., 2008; Zhekov & Tomov, 2019), SS73 17 (CD–57 3057; Smith et al., 2008; Eze et al., 2010), and MWC 560 (Stute & Sahai, 2009; Lucy et al., 2020). The discovery of this particular group presents a challenge to our knowledge about accretion physics in white dwarfs due to the strong hard X-ray emission (Luna & Sokoloski, 2007; Kennea et al., 2009). Recently, Toalá (2024) proposed a disk-like model to explain the X-ray properties of symbiotics, where -type sources have an accretion disk near the edge. In addition, radiative transfer simulations of X-ray photons by Toalá (2024) implied that the -type group is likely related to low-accreting degenerate cores with high-temperature plasma ( keV) within the boundary layer between the inner edge of the accretion disk and the white dwarf surface.
@@
Some hard X-ray-emitting symbiotics seem to produce distinctive soft and hard thermal plasma emissions: CH Cyg with plasma temperatures of 0.2, 0.7, and 7.3 keV (Ezuka et al., 1998), SS73 17 with temperatures of and keV (Eze et al., 2010), and MWC 560 showing thermal emissions with temperatures of and keV (Stute & Sahai, 2009). Symbiotic stars with distinctive soft and hard X-ray thermal components are referred to as the -type group (Luna et al., 2013). Previous studies have also identified the presence of jets in some of them, namely CH Cyg (Galloway & Sokoloski, 2004; Karovska et al., 2007, 2010) and MWC 560 (Tomov et al., 1992; Schmid et al., 2001; Lucy et al., 2018). The soft thermal emission found in these systems might have a potential link to the shock region created by either a jet or the interaction of a wind colliding with the surrounding material (Stute & Sahai, 2009). According to Toalá (2024), the soft X-ray emission of a two-temperature plasma model can be obscured by the disk, so extended emission from jets, colliding winds, and/or hot bubbles is likely responsible for the soft component in sources. The soft X-ray component could originate from colliding stellar winds (CSWs), though a different mechanism such as accretion may also be responsible for the X-ray emission seen in some epochs in the -class symbiotic star AG Peg (Zhekov & Tomov, 2016).
@@
The X-ray variability of RT Cru recorded with the Chandra telescope has been recently investigated using hardness ratio analysis and spectral modeling (Danehkar et al., 2021). To further evaluate the nature of the X-ray variability in this -type symbiotic star and the soft thermal plasma emission suggested by Danehkar et al. (2021), we conduct further timing analyses on the archival data taken with XMM-Newton, Suzaku, and NuSTAR, in addition to comprehensive eigenvector-based multivariate analyses of the historical X-ray data of the Chandra and other telescopes. Section 2 describes our reduction of time-sliced data required for implementing principal component analysis (PCA). In Section 3, we investigate the X-ray light curves and hardness ratios. Section 4 presents our principal components determined with PCA, as well as simulated PCA spectra, and is followed by discussions in Section 5 and a conclusion in Section 6.
@@ 2 Data Reduction for PCA
RT Cru was observed using the X-ray Imaging Spectrometer (XIS; Koyama et al., 2007) and the Hard X-ray Detector (HXD; Takahashi et al., 2007) on board Suzaku (Mitsuda et al., 2007) in 2007 and 2012, the EPIC-pn instrument (Strüder et al., 2001) aboard the XMM-Newton telescope (Jansen et al., 2001) in 2019, the ‘A’ and ‘B’ focal plane modules (FPM) on the Nuclear Spectroscopic Telescope Array (NuSTAR) satellite (Harrison et al., 2013) in 2016, and the Low Energy Transmission Grating (LETG; Brinkman et al., 2000) on the High Resolution Camera Spectrometer (HRC-S; Murray et al., 2000) of Chandra X-ray Observatory (CXO; Weisskopf et al., 2000, 2002) in 2015. The observations are summarized in Table 1, which includes the instrument and its configuration, observation identification number, start and end times (UTC), exposure time (ks), and total counts and count rate (count s-1) of each observation. To implement eigenvector-based multivariate statistical analysis such as PCA of a variable X-ray source, it is necessary to split each dataset into a time series of spectral data at fixed intervals of 10 ks.
@@
We downloaded the Chandra HRC-S/LETG event data from the Chandra data archive and reprocessed them using the chandra_repro tool from the ciao package (version 4.15; Fruscione et al., 2006) together with the Chandra CALDB data (version 4.10.2).111The Chandra dataset is contained in https://doi.org/10.25574/cdc.201 (catalog doi:10.25574/cdc.201). The time-segmented event files were produced by applying the ciao operations dmcopy and dmappend on the 2nd level events. The low-energy grating (LEG) spectra, together with their respective redistribution and response data, were generated by applying the ciao programs tgextract and mktgresp to the time-sliced event files. The dmtype2split tool, part of the ciao software, was utilized to segregate distinct positive and negative orders, while the application tg_bkg was used to create the background spectra.
@@
The XMM-Newton EPIC-pn data were obtained from XMM Science Archive and reduced with the science analysis software (sas v 20.0.0; Gabriel et al., 2004) and the calibration files (XMM-CCF-REL-391). The use of the sas tool epproc led to the generation of processed event files. These events were subsequently employed to produce new event files, which were stacked at intervals of 10 ks with the sas program evselect. We removed the events of flaring particle background, which were identified using count rates of c s-1 in the single-pixel (PATTERN 0) light curves binned at 100 sec within the 10–12 keV energy range. Furthermore, the time-filtered event files exclusively included single and double patterned events (PATTERN 4) within the relevant pulse-invariant (PI) channel range (200 PI 15000), disregarding defective pixels (FLAG 0). The procedure especget was utilized to generate a set of source spectra by applying it to the time-sliced event data. The spectra of the source were taken from a circular region with a radius of centered on the brightest peak of the source. The background spectra were created using a circle of the same size on the same chip, but without any sources included.
@@
The Suzaku data acquired using the XIS and HXD-PIN were retrieved and processed using the ftool program aepipeline from the HEAsoft software (v 6.31.1) and the calibration data (XIS: 2018-10-10 and HXD: 2011-09-13). The XIS provided energy coverage ranging from 0.2 to 12 keV, while the HXD supplied a band pass of 10-70 keV using PIN diodes. Various Good Time Interval (GTI) files at our chosen time interval were created using the ftool application maketime. The GTI files were employed in the multipurpose tool xselect to generate a set of time-segmented source spectra. The source spectrum was acquired by extracting data from a circular region with a radius of centered on the emission peak of the source, whereas the background was chosen from a nearby circular region of the same radius excluding any sources. We should specify that XIS1 was the only back-illuminated (BI) detector aboard Suzaku, whereas the other XIS devices were front-illuminated (FI). Typically, BI detectors are expected to offer superior quantum efficiency to FI detectors for an identical depletion depth (Lesser & Iyer, 1998). The program hxdpinxbpi was employed to generate the time-sliced spectra of HXD-PIN data using the GTI files and the ”tuned” non-X-ray PIN background released by the Suzaku team.222https://heasarc.gsfc.nasa.gov/FTP/suzaku/data/background/
@@
The NuSTAR data taken with the FPM ‘A’ and ‘B’ were downloaded from the HEASARC archive and reduced using the tool nupipeline from NuSTARDAS (v 2.1.2) and the relevant calibration files (8 Feb 2023). The ftool program maketime was used to create the GTI tables at intervals of 10 ks, which were then applied to the calibrated event files via the application nuproducts resulting in the time-stacked spectra. The source was extracted from a circular aperture with a radius of , whereas the background was from a location on the same chip that was devoid of any sources.
@@ 3 Time Series Analysis
The first two equations are identical to those employed by Prestwich et al. (2003) for the classification of X-ray sources in the Local Group galaxies. The hardness ratio diagrams are created by plotting the hardness ratios against the entire bands, and assist in the detection of spectral transitions associated with accretion or obscuration caused by absorbing material. Hardness ratio analysis has been employed to characterize various astronomical objects, including extragalactic X-ray sources (e.g., Hong et al., 2004; Plucinsky et al., 2008), quasars (Danehkar et al., 2018; Boissay-Malaquin et al., 2019), and X-ray binaries (Sreehari & Nandi, 2021).
In Figure 1, we present the background-subtracted light curves of RT Cru made with the XMM-Newton, Suzaku, and NuSTAR observations binned at intervals of 600 sec in the broad (; 0.4–10 keV), hard (), and extreme hard () bands. The figure also shows the corresponding hardness ratios, namely and for XMM-Newton (pn) and Suzaku (XIS1), and for NuSTAR and Suzaku (XIS1+PIN). The uncertainties in the light curves and hardness ratios were determined using the Bayesian Estimator for Hardness Ratios (BEHR; Park et al., 2006). Hour-scale variations are evident in all the time series. In particular, the source got harder in 2012 February compared to 2007 July according to the time series of the Suzaku XIS1 data, whereas the source brightness () in 2012 is lower than that in 2007. Moreover, the and hardness ratios involving the extreme hard band () were stronger in 2012 February.
Figure 2 shows hardness ratio diagrams that illustrate the relationship between the hardness ratios and the broad bands ( and ). The hardness diagrams of the XMM-Newton observation do not depict a hardness transition similar to what was seen with the Chandra data (Danehkar et al., 2021). However, the and diagrams of the two Suzaku observations indicate that the source experienced a transition from the high/soft to low/hard spectral states between 2007 and 2012. Moreover, the Suzaku diagrams also revealed that the source was stronger in the soft (0.4–1.1 keV) and hard (2.6–10 keV) bands – lower and higher at the same time – in 2012 February by comparison with those in 2007 July. Although we do not see any special pattern in the hardness ratio diagrams of NuSTAR, the and diagrams of the Suzaku light curves imply a higher average of the extreme hardness () in 2012 February.
@@ 3.1 Statistical Tests of X-ray Variability
where and are the standard deviation and the mean of data points, respectively, is the index of each data point in the time series, and is the total number of data points. This coefficient describes the extent of variations with respect to the mean, so its higher values are associated with stronger variability if we compare the same parameter. However, this coefficient is not suitable for comparison between dissimilar parameters and cannot evaluate the intrinsic nature of variations. The ratio of the standard deviation of errors to that of data points () may help us see whether the variability is intrinsic to the observation. Again, this ratio is suitable for evaluating values for the identical parameter. Both and are unable to characterize detailed features of consecutive series. The normalized consecutive number, so-called Con, which was first used by Wozniak (2000), may be used to quantify the changes at a number of consecutive points. To evaluate continuity in light curves, we define Con as the number of three consecutive points greater or lower than normalized by . For random fluctuations, Con has a value of and is equal to 0 for constant consecutive series.
@@
For a normal distribution, the mean von Neumann ratio is expected to be (Young, 1941). To assess the presence of autocorrelation, we estimated the confidence intervals of with the significance level of for a normal probability distribution. A von Neumann ratio within the confidence levels of implies no autocorrelation, whereas values outside the aforementioned confidence levels toward 0 and 4 correspond to positive and negative autocorrelation in successive data points, respectively.
@@
Table 2 summarizes the results of our statistical analysis of the X-ray variability in RT Cru obtained with different methods. It can be seen that and of , HR1, and HR2, are higher in the XMM-Newton data than those in the Suzaku observations. This implies that variations are higher in the XMM-Newton light curves, but with larger uncertainties. We also notice a higher in the HR4 ratio of the NuSTAR data, which may be an indication of scattered fluctuations in this ratio. However, is not a suitable tool for quantifying the variability in a time sequence of data. The Con number may be able to better distinguish variations in a consecutive sequence. The values of Con indicate consecutive changes in in the XMM-Newton and combined Suzaku/XIS1 observations, HR2 in the first Suzaku/XIS1 observation and two combined Suzaku/XIS1 observations, while others with Con may have random variations. Nevertheless, our Con statistical method is based only on three consecutive data points, so it is unable to obtain a broader picture of variability in the entire sequence of data points.
@@
In the case of Suzaku/XIS1, the von Neumann ratio demonstrates positive autocorrelation in the time series of the broad band () and hardness ratios of the first, second, and combined observations, apart from HR1 in the second observation. However, the -values of the normality tests suggest that there are not normal distributions in HR1 of the first XIS1 observation with marginally significant statistics and the second XIS1 observation with significant statistics. Furthermore, the A–D and S–W tests of , HR1, and HR2 in the two mixed Suzaku/XIS1 observations definitely reject the hypothesis of normality, which is obvious in the movement pattern seen in the hardness diagrams in Fig. 2. Nevertheless, we see a non-normal distribution only in the HR3 ratio of the Suzaku/PIN observations, but not and HR4. The von Neumann statistical analysis of the Suzaku/PIN data also depicts positive autocorrelation on the broad band of the first and combined observations, and on HR3 and HR4 of the mixed multi-epoch data, which are consistent with the spectral transition occurring between 2007 and 2012.
The broad band () and the hardness ratios (HR3 and HR4) of the NuSTAR data exhibit non-random distributions according to the -values obtained from A–D and S–W tests. However, von Neumann’s mean squared successive difference depicts a positive autocorrelation only in the broad band. The hardness diagrams in Fig. 2 seem to not depict any obvious hardness transition in the NuSTAR data, though our A–D and S–W normality tests apparently suggest it, which is not supported by the von Neumann statistics. However, all of the normality tests and the von Neumann ratio indicate that there are statistically significant fluctuations in the NuSTAR broad band.
In summary, our statistical analysis of the variability shows that RT Cru underwent flux variations over the broad band of the XMM-Newton data based on and the A–D/S–W test, the first and second Suzaku/XIS1 data, the first and combined Suzaku/PIN data, the NuSTAR data according to the von Neumann results, and the mixed first-second Suzaku/XIS1 data and the NuSTAR data based on the A–D/S–W test. Moreover, our results suggest that the X-ray source experienced a long-term spectral transition between the first and second Suzaku observations based on the normality tests of HR1, HR2, and HR3, as well as some short-term hardness fluctuations in HR1 of the second Suzaku/XIS1 data and HR3 of the NuSTAR data.